3 suppose ‘X’ is normally distributed with mean 20 and variance 8 i.e X-NID


Find


A P(25≤X≤30)


B P(X≥25)


4 Given the following discrete probability destruction.


X 1 5 7 9


F(x) 1/6 2/6 2/6 1/6



Find


A the expected value of x E(X)


The variance of x V(X)


6 the probabilities that Abebe passes microeconomics is 2/3 and the probability that he passed statics 4/9. If the probability of passing both courses is 1/4 ,what is the probability that Abebe will pass at lease one of these courses

1
Expert's answer
2022-06-03T13:18:12-0400

3.

A.


"P(25\le X\le30)=P(Z\le\dfrac{30-20}{\sqrt{8}})"

"-P(Z<\dfrac{25-20}{\sqrt{8}})\approx P(Z\le3.5355)"

"-P(Z<1.7678)\approx0.0383"

B.


"P( X\ge25)=1-P(Z<\dfrac{25-20}{\sqrt{8}})"

"\approx1-P(Z<1.7678)\approx0.0385"

4.

A.


"E(X)=\dfrac{1}{6}(1)+\dfrac{2}{6}(5)+\dfrac{2}{6}(7)+\dfrac{1}{6}(9)"


"=\dfrac{17}{3}"

B.


"E(X^2)=\dfrac{1}{6}(1)^2+\dfrac{2}{6}(5)^2+\dfrac{2}{6}(7)^2+\dfrac{1}{6}(9)^2=\dfrac{115}{3}"

"Var(X)=E(X^2)-(E(X))^2=\dfrac{115}{3}
-(\dfrac{17}{3})^2"

"=\dfrac{56}{9}"

6.


"P(M\cup S)=P(M)+P(S)-P(M\cap S)"

"=\dfrac{2}{3}+\dfrac{4}{9}-\dfrac{1}{4}=\dfrac{31}{36}"